# tda-segmentor: A tool to extract and analyze local structure and porosity features in porous materials

### v1

@@ 1 Introduction
Porous materials of various chemical compositions and pore size-scales are both commonly observed in nature and applied in technology, and have long attracted researcher’s attention. For example, nanoporous materials such as zeolites, natural clays and metal organic frameworks (MOF) are finding applications in the industry as chemical catalysts KONNERTH2020213319 , membranes or as adsorbents Auerbach2003 . Porous metal structures are investigated in the context of lightweight structural materials for energy-efficient transport PATEL201820391 . Numerous others are investigated in the context of battery electrodes EGOROV2020201 , bio-materials zadpoor2019additively , geological deposits and others.
@@
Computational approaches have been playing a crucial role in understanding the structure-property relationships and aiding the design. For example, molecular simulations can be use to reliably predict adsorption and diffusion of guest species Smit2008 ; Krishna2007 . Similarly, in porous metals such as gold, simulations can shed light onto relationships between the topology of the structure and/or its pore network and the mechanical properties of the material. For instance, finite element simulations have shown that for two porous structures with same solid fractions but different pore topologies (gyroids vs spinodal-like) yield at different compressive stresses (Mangipudi2016, ). In addition through molecular dynamics simulations, compression of nanoporous structures with the same solid fraction and ligament diameter but different topology are found to have different mechanical responses (Mathesan2021, ; Mathesan2020, ).
@@
Furthermore, in recent years a wide accessibility of high-throughput computing has allowed studies of larger sets of structures. Machine learning-based approaches have become to emerge, i.e. by training on a small set of simulated data, ML regressors can predict the material property just from the structure features (descriptors). For guest adsorption for example, porosity descriptors such as accessible surface area (ASA), largest cavity diameter (LCD) can readily generate feature vectors which can act as an input for the machine learning algorithms Simon2015 . However such pore descriptors only work well at high pressures where the guest molecule adsorbs in the entire void space, but for low pressures they tend to be localized in the strongly binding regions of the material’s pore (Lin2012, ). Hence there is a need to generate robust descriptors that is also able to include extremely local pore morphology of the structures too.
@@
Topological Data Analysis (TDA) has already developed into a mature field that is capable of obtaining insight from large datasets by looking at their ‘shape’ and how it can reveal important information of the data. In particular, persistent homology (edelsbrunner2022, ) can explain how certain features of the data persist across multiple-scales. Such methods are already popular in network science (Aktas2019, ) and have also been recently used for applications in material science. In (Lee2017, ), persistent homology is used to formulate robust porosity descriptors and they show that two structures with similar pore morphology have so-called similar persistent diagrams. In (Krishnapriyan2020, ), such persistent diagrams have been converted to persistent images which can then act as feature vectors for machine learning algorithms.
@@
In this work, we develop a high-throughput tool that borrows from these ideas in Topological Data Analysis. In particular, from Morse theory, which in algebraic topology is used to analyze the topology of a manifold by studying differentiable functions on that manifold, cell-complexes of the manifold can be formed, which can be used in decomposing the manifold into different unique regions. We present here a tool written in C++ that is built on the topology toolkit library (TTK) (ttk, ), which have efficient algorithms for the topological analysis of scalar functions, and we build on these algorithms to analyze our data. The rest of the article is organized as follows: In section 2, we first present the TDA method to generate the segments. We discuss the different input fields that can be used for analysis, the methodology based on persistence used to filter noise and finally generate unique segments. In section 3 we apply it to some example zeolites and nanoporous gold structures to generate segments of either the pore or the structure. Further, using the segmentation information, we also demonstrate that we can replace the entire structure by a much simpler graph representation, and finally in section 4, we discuss some applications from the generated segments related to guest adsorption and pore-similarity. We conclude with some remarks in section 5.
@@ 2 Methods
Topological Data Analysis (TDA) is premised on the idea that the shape of data sets contains relevant information. Herein, we first present some of the mathematical preliminaries of topological analysis that is based on persistent homology and we refer to (Milnor, ; Forman2002, ; edelsbrunner2022, ) for an in-depth analysis of the theory. TDA has seen successful applications in a variety of fields of science such as combustion (Bremer, ; Gyulassy2014, ), astrophysics (Shivashankar, ) and materials science (Gyulassy2016, ; Lee2017, ). The following subsections describe our implementation and the input data.
@@ 2.1 Mathematical overview
Let us consider a manifold $\mathcal{M}$ and a piece wise linear scalar field $f:\mathcal{M}\to\mathbb{R}$, where $d$ is the dimension of the manifold $\mathcal{M}$. While the theoretical framework can be extended to any dimension $d$, we restrict ourselves here to $d=3$, which is the three dimensional Cartesian space. The function $f$ has values at the vertices $\mathcal{M}^{0}$ of $\mathcal{M}$ and for higher dimension simplices, such as edges, faces or the volume within the faces, the function $f$ is linearly interpolated. The first step is to evaluate the critical points of the scalar function $f$. For smooth functions, the critical points are simply the points where the gradient of the scalar field, $\nabla f$ vanishes, while in the discrete setting, the critical points are evaluated slightly differently (see (ttk, )). Once the critical points are evaluated, they are given indices ($\mathcal{I}$) with $\mathcal{I}=0$ for minima, $\mathcal{I}=1$ for 1-saddles, $\mathcal{I}=2$ for 2-saddles and $\mathcal{I}=3$ for maxima.
@@
Figure 1: Morse-Smale segmentation example of a noisy terrain taken from (ttk, ): (a) shows the noisy surface $f$ and (b) shows all the critical points of the noisy surface from which we plot the persistent diagram $\mathcal{D}(f)$ in (e) and persistence curve $\mathcal{C}(f)$ in (d). (c) plots the persistence diagram after the application of a persistent threshold (red line in (d)). Once the noisy features are discarded, we get the unique Morse-Smale segmentation (f) of the terrain. The same information can also be used to segment the surface into unique ascending and descending manifolds as shown in (g) and (h) respectively.
@@
Let us consider an example from the documentation of the TTK library where the scalar field $f$ is a discrete wavy terrain as shown in Fig. 1 (a) where there are clear hills and valleys with some noise. If the critical points for this surface is calculated as shown by the spheres in Fig. 1(b), critical points are also located in all the noisy features of the terrain. The first step in TDA is to look at the distribution of critical points of $f$ through the so-called persistence curve, $\mathcal{C}(f)$, and the persistence diagram, $\mathcal{D}(f)$. Critical points can be arranged in a set of pairs, such that each critical point appears in only one pair $(c_{i},c_{j})$ with $f(c_{i})<f(c_{j})$ and $\mathcal{I}(c_{i})=\mathcal{I}(c_{j})-1$ (edelsbrunner2022, ). If one plots on the X-axis, $f(c_{i})$ and on the Y-axis a bar with ordinates $f(c_{i})$ and $f(c_{j})$, then we obtain the persistence diagram, $\mathcal{D}(f)$, as shown in Fig. 1 (e) and we call $p=f(c_{j})-f(c_{i})$ as the persistence. The noisy features of the terrain are less persistent and hence populate as bars very close to the diagonal. A persistence threshold can then be applied on this diagram (see Fig. 1(c)), to remove the noisy features, and evaluate the more significant features as shown by spheres in Fig. 1 (f). Alternatively, the number of critical pairs $(c_{i},c_{j})$ can also be plotted with the persistence resulting in the persistence curve, $\mathcal{C}(f)$, as shown in Fig. 1 (d). Critical pairs at a low value of persistence that correspond to the noisy features disappear quickly, and separate from the significant features through a horizontal plateau as shown by the red straight line in Fig. 1(d). Such a horizontal plateau is quite typical in analyzing scalar data as it clearly separates scales from smaller to larger features and can act as a good guide to decide the persistence threshold for applications.
@@ 2.2 Implementation
Our tda-segmentor tool that can be used to apply the aforementioned ideas on nanoporous structures such as, zeolites, MOF’s or metallic nano-pillars. The code is open-source, developed in C++ and is published on GitHub. The algorithms for topological analysis are taken from the Topology Toolkit (TTK) (ttk, ), which is a software platform that is easily accessible to the end users due to a tight integration with Paraview. Moreover, the library comes with a variety of bindings in Python and VTK/C++ for fast prototyping or even through direct C++ without any VTK dependency. In our code, we take advantage of the VTK/C++ dependency as it results in much cleaner, efficient and better code.
@@ 2.3.1 Distance grids
As various chemical properties such as adsorption of a molecule directly depends on the pore morphology, the distance of a point to the material internal surface provides useful connections between structure and property. Additionally, we can use the sign of the distance function to extend the same algorithmic framework to study not only porosity but also the corresponding material structures. Specifically, we define the distance grid to be to be the least distance to the surface of the structure and negative within the structure while positive in the pore (see Fig. 3(a) for the zeolite FAU and Fig. 7(b) for gold nano-pillar structures). For zeolites, the distance grids can be evaluated using Zeo++(Willems2012, ) that accepts common crystal structure input formats such as .cif and .cssr.
@@ 2.3.2 Energy grids
When a guest molecule enters the porous materials, it faces repulsive and attractive forces resulting in a complex energy landscape that dictates, together with other parameters such as temperature, how the molecule diffuses and/or adsorbs within the pore network. The guest molecule binding sites correspond to the minima of the energy landscape. Interaction energy grids (to be referred to as simply energy grids) can be evaluated from an existing open source repository PorousMaterials.jl(PorousMaterials, ) that may use fairly standard Lennard-Jones potential based on the universal force field. An example for an energy grid is shown in Fig. 3 (b), for a methane (CH4) molecule into a LTA zeolite framework.

### v2

@@ 1 Introduction
Porous materials of various chemical compositions and pore size-scales are both commonly observed in nature and applied in technology, and have long attracted researcher’s attention. For example, nanoporous materials such as zeolites, natural clays and metal organic frameworks (MOF) are finding applications in the industry as chemical catalysts [1], membranes or as adsorbents [2]. Porous metal structures are investigated in the context of lightweight structural materials for energy-efficient transport [3]. Numerous others are investigated in the context of battery electrodes [4], bio-materials [5], geological deposits and others.
@@
Computational approaches have been playing a crucial role in understanding the structure-property relationships and aiding the design. For example, molecular simulations can be use to reliably predict adsorption and diffusion of guest species [6, 7] . Similarly, in porous metals such as gold, simulations can shed light onto relationships between the topology of the structure and/or its pore network and the mechanical properties of the material. For instance, finite element simulations have shown that for two porous structures with same solid fractions but different pore topologies (gyroids vs spinodal-like) yield at different compressive stresses [8]. In addition through molecular dynamics simulations, compression of nanoporous structures with the same solid fraction and ligament diameter but different topology are found to have different mechanical responses [9, 10].
@@
Furthermore, in recent years a wide accessibility of high-throughput computing has allowed studies of larger sets of structures. Machine learning-based approaches have become to emerge, i.e. by training on a small set of simulated data, ML regressors can predict the material property just from the structure features (descriptors). For guest adsorption for example, porosity descriptors such as accessible surface area (ASA), largest cavity diameter (LCD) can readily generate feature vectors which can act as an input for the machine learning algorithms [11]. However such pore descriptors only work well at high pressures where the guest molecule adsorbs in the entire void space, but for low pressures they tend to be localized in the strongly binding regions of the material’s pore [12]. Hence there is a need to generate robust descriptors that is also able to include extremely local pore morphology of the structures too.
@@
Topological Data Analysis (TDA) has already developed into a mature field that is capable of obtaining insight from large datasets by looking at their ‘shape’ and how it can reveal important information of the data. In particular, persistent homology [13] can explain how certain features of the data persist across multiple-scales. Such methods are already popular in network science [14] and have also been recently used for applications in material science. In [15], persistent homology is used to formulate robust porosity descriptors and they show that two structures with similar pore morphology have so-called similar persistent diagrams. In [16], such persistent diagrams have been converted to persistent images which can then act as feature vectors for machine learning algorithms.
@@
In this work, we develop a high-throughput tool that borrows from these ideas in Topological Data Analysis. In particular, from Morse theory, which in algebraic topology is used to analyze the topology of a manifold by studying differentiable functions on that manifold, cell-complexes of the manifold can be formed, which can be used in decomposing the manifold into different unique regions. We present here a tool written in C++ that is built on the topology toolkit library (TTK) [17], which have efficient algorithms for the topological analysis of scalar functions, and we build on these algorithms to analyze our data. The rest of the article is organized as follows: In section 2, we first present the TDA method to generate the segments. We discuss the different input fields that can be used for analysis, the methodology based on persistence used to filter noise and finally generate unique segments. In section 3 we apply it to some example zeolites and nanoporous gold structures to generate segments of either the pore or the structure. Further, using the segmentation information, we also demonstrate that we can replace the entire structure by a much simpler graph representation, and finally in section 4, we discuss some applications from the generated segments related to guest adsorption and pore-similarity. We conclude with some remarks in section 5.
@@ 2 Methods
Topological Data Analysis (TDA) is premised on the idea that the shape of data sets contains relevant information. Herein, we first present some of the mathematical preliminaries of topological analysis that is based on persistent homology and we refer to [18, 19, 13] for an in-depth analysis of the theory. TDA has seen successful applications in a variety of fields of science such as combustion [20, 21], astrophysics [22] and materials science [23, 15]. The following subsections describe our implementation and the input data.
@@ 2.1 Mathematical overview
Let us consider a manifold $\mathcal{M}$ and a piece wise linear scalar field $f:\mathcal{M}\to\mathbb{R}$, where $d$ is the dimension of the manifold $\mathcal{M}$. While the theoretical framework can be extended to any dimension $d$, we restrict ourselves here to $d=3$, which is the three dimensional Cartesian space. The function $f$ has values at the vertices $\mathcal{M}^{0}$ of $\mathcal{M}$ and for higher dimension simplices, such as edges, faces or the volume within the faces, the function $f$ is linearly interpolated. The first step is to evaluate the critical points of the scalar function $f$. For smooth functions, the critical points are simply the points where the gradient of the scalar field, $\nabla f$ vanishes, while in the discrete setting, the critical points are evaluated slightly differently (see [17]). Once the critical points are evaluated, they are given indices ($\mathcal{I}$) with $\mathcal{I}=0$ for minima, $\mathcal{I}=1$ for 1-saddles, $\mathcal{I}=2$ for 2-saddles and $\mathcal{I}=3$ for maxima.
@@
Figure 1: Morse-Smale segmentation example of a noisy terrain taken from [17]: (a) shows the noisy surface $f$ and (b) shows all the critical points of the noisy surface from which we plot the persistent diagram $\mathcal{D}(f)$ in (e) and persistence curve $\mathcal{C}(f)$ in (d). (c) plots the persistence diagram after the application of a persistent threshold (red line in (d)). Once the noisy features are discarded, we get the unique Morse-Smale segmentation (f) of the terrain. The same information can also be used to segment the surface into unique ascending and descending manifolds as shown in (g) and (h) respectively.
@@
Let us consider an example from the documentation of the TTK library where the scalar field $f$ is a discrete wavy terrain as shown in Fig. 1 (a) where there are clear hills and valleys with some noise. If the critical points for this surface is calculated as shown by the spheres in Fig. 1(b), critical points are also located in all the noisy features of the terrain. The first step in TDA is to look at the distribution of critical points of $f$ through the so-called persistence curve, $\mathcal{C}(f)$, and the persistence diagram, $\mathcal{D}(f)$. Critical points can be arranged in a set of pairs, such that each critical point appears in only one pair $(c_{i},c_{j})$ with $f(c_{i})<f(c_{j})$ and $\mathcal{I}(c_{i})=\mathcal{I}(c_{j})-1$ [13]. If one plots on the X-axis, $f(c_{i})$ and on the Y-axis a bar with ordinates $f(c_{i})$ and $f(c_{j})$, then we obtain the persistence diagram, $\mathcal{D}(f)$, as shown in Fig. 1 (e) and we call $p=f(c_{j})-f(c_{i})$ as the persistence. The noisy features of the terrain are less persistent and hence populate as bars very close to the diagonal. A persistence threshold can then be applied on this diagram (see Fig. 1(c)), to remove the noisy features, and evaluate the more significant features as shown by spheres in Fig. 1 (f). Alternatively, the number of critical pairs $(c_{i},c_{j})$ can also be plotted with the persistence resulting in the persistence curve, $\mathcal{C}(f)$, as shown in Fig. 1 (d). Critical pairs at a low value of persistence that correspond to the noisy features disappear quickly, and separate from the significant features through a horizontal plateau as shown by the red straight line in Fig. 1(d). Such a horizontal plateau is quite typical in analyzing scalar data as it clearly separates scales from smaller to larger features and can act as a good guide to decide the persistence threshold for applications.
@@ 2.2 Implementation
Our tda-segmentor tool that can be used to apply the aforementioned ideas on nanoporous structures such as, zeolites, MOF’s or metallic nano-pillars. The code is open-source, developed in C++ and is published on GitHub. The algorithms for topological analysis are taken from the Topology Toolkit (TTK) [17], which is a software platform that is easily accessible to the end users due to a tight integration with Paraview. Moreover, the library comes with a variety of bindings in Python and VTK/C++ for fast prototyping or even through direct C++ without any VTK dependency. In our code, we take advantage of the VTK/C++ dependency as it results in much cleaner, efficient and better code.
@@ 2.3.1 Distance grids
As various chemical properties such as adsorption of a molecule directly depends on the pore morphology, the distance of a point to the material internal surface provides useful connections between structure and property. Additionally, we can use the sign of the distance function to extend the same algorithmic framework to study not only porosity but also the corresponding material structures. Specifically, we define the distance grid to be to be the least distance to the surface of the structure and negative within the structure while positive in the pore (see Fig. 3(a) for the zeolite FAU and Fig. 7(b) for gold nano-pillar structures). For zeolites, the distance grids can be evaluated using Zeo++[24] that accepts common crystal structure input formats such as .cif and .cssr.
@@ 2.3.2 Energy grids
When a guest molecule enters the porous materials, it faces repulsive and attractive forces resulting in a complex energy landscape that dictates, together with other parameters such as temperature, how the molecule diffuses and/or adsorbs within the pore network. The guest molecule binding sites correspond to the minima of the energy landscape. Interaction energy grids (to be referred to as simply energy grids) can be evaluated from an existing open source repository PorousMaterials.jl[25] that may use fairly standard Lennard-Jones potential based on the universal force field. An example for an energy grid is shown in Fig. 3 (b), for a methane (CH4) molecule into a LTA zeolite framework.